# Scale Reliant Inference

### v4

@@ Abstract
Scientific fields such as genomics, ecology, and political science often collect multivariate count data. In these fields, the data are often sufficiently noisy such that inferences regarding the total size of the measured systems have substantial uncertainty. This uncertainty can hinder downstream analyses, such as differential analysis in case-control studies. There have historically been two approaches to this problem: one considers the data as compositional and the other as counts that can be normalized. In this article, we use the framework of partially identified models to rigorously study the types of scientific questions (estimands) that can be answered (estimated) using these data. We prove that satisfying Frequentist inferential criteria is impossible for many estimation problems. In contrast, we find that the criteria for Bayesian inference can be satisfied, yet it requires a particular type of model called a Bayesian partially identified model. We introduce Scale Simulation Random Variables as a flexible and computationally efficient form of Bayesian partially identified models for analyzing these data. We use simulations and data analysis to validate our theory.
Keywords: Multivariate Analysis, Identifiability, Uncertainty Quantification, Sequence Count Data
@@ 1 Introduction
Many scientific fields collect noisy multivariate measurements where the scale of the measurements (i.e., total measured amount) is arbitrary and does not represent the scale of the underlying system (i.e., true total amount). These data occur frequently in genomics (Lovell et al.,, 2011; Gloor et al.,, 2017), macro-ecology (Murphy et al.,, 2016; Jackson,, 1997), epidemiology (Silverman et al.,, 2020), demography (Rossi et al.,, 1987), and political polling (Topîrceanu and Precup,, 2020). Consider forecasting a national election based on a collection of polls conducted in distinct sub-national regions. The scale of the polls may be incomparable such that a single vote in one poll represents a different number of total votes than a single vote in a different poll. The observed counts could be normalized for downstream analysis using survey weighting if accurate information on the size of the voter roll is available for each region. However, the problem is substantially more complicated if such data is lacking or subject to errors. Similar challenges are faced when studying gut microbiota, where fecal samples are sequenced to quantify the abundance of different microbes in each subject’s gut. Prior work has established that the measurement scale (e.g., the total number of sequenced DNA molecules) is driven by technical factors: it is unrepresentative of meaningful biological variation in total microbial load (Lovell et al.,, 2011; Props et al.,, 2017; Vandeputte et al.,, 2017; Paulson et al.,, 2013). Statisticians often agree that these issues of arbitrary data scale can confound analyses (Topîrceanu and Precup,, 2020). For example, in studying microbiota, it is well known that methods used to account for this arbitrary scale can dominate study conclusions (Roche and Mukherjee,, 2022; Weiss et al.,, 2017). However, there are fundamental disagreements regarding how to address these challenges and what types of analyses can be rigorously performed (Gloor et al.,, 2017; Lovell et al.,, 2015, 2011; Gloor et al.,, 2017). This paper addresses the question: when does the arbitrary scale of the data limit the types of analyses that can be accurately carried out and which methods are best suited for mitigating undesirable inference properties such as bias and Type-I error?
Some authors call these data compositional (Gloor et al.,, 2017). This term reflects the idea that if the data scale is arbitrary, then the data only contains information about the system’s composition (i.e., proportions). According to the Scale Invariance Principle of Compositional Data Analysis (CoDA), if the observed data are compositional, then analyses must be scale invariant: they must focus solely on the relative frequency of the counts (Greenacre et al.,, 2023; Pawlowsky-Glahn et al.,, 2015). For example, within the study of human microbiota, proponents of the CoDA perspective have argued that inter-taxa correlations cannot be estimated rigorously as they are a function of microbial load and cannot be expressed in terms of microbial compositions alone (Lovell et al.,, 2015). Instead, CoDA proponents have suggested that authors focus their attention on scale invariant statistics such as proportionality (Lovell et al.,, 2015). In short, the field of CoDA develops tools to answer scale invariant research questions and often dismisses scale reliant questions as unanswerable (Aitchison,, 1982; Lovell et al.,, 2015, 2011).
CoDA’s scale invariant perspective is rigorous, but the rigor comes with constraints on the scientific question. While constraining ourselves to scale invariant questions seems logical when we cannot estimate the system scale accurately, it is sub-optimal in less extreme settings. In the setting of polling, additional measurements of scale (e.g., the size of the voter roll) could be used to overcome these limits and rigorously answer scale reliant questions. Alternatively, even if we lack scale measurements, we may have prior knowledge that allows us to quantify our uncertainty about the system scale. This paper formally defines the conditions under which we can move beyond CoDA’s insistence on scale invariance and answer scale reliant questions using only incomplete information about the system scale.
While this paper is the first to formalize the conditions under which scale reliant questions can be answered using data where the scale/total counts are noisy or missing, many models, algorithms, and methods have been proposed to address common scale reliant inferential tasks such as correlation analysis or differential analysis (Lovell et al.,, 2015; Fernandes et al.,, 2014; Love et al.,, 2014). These include data normalizations (Weiss et al.,, 2017), categorical data models with identification restrictions (Rabe-Hesketh and Skrondal,, 2001), or external measurements of scale (Gibson and Gerber,, 2018; Williamson et al.,, 2019), as well as a wide variety of methods that assume sparsity in the underlying system (Grantham et al.,, 2020). There are also many other approaches, and we cannot review them all. Instead, motivated by recent debates within the genomics literature (e.g.,  Gloor et al., (2017)), we bias our presentation towards examples from that field to make the theory in this paper concrete.
The main contributions of this paper include the following. Using the concept of partial identifiability (Gustafson,, 2015), we state the conditions under which the estimation of estimands of scientific interest, such as correlations between variables or differential analysis between case versus control, are possible and reliable. We show that Frequentist procedures do not provide reliable estimation of these estimands. Then, we demonstrate that a particular class of Bayesian models called Bayesian Partially Identified Models (Bayesian PIMs) can result in rigorous inference. We also provide a simulation framework called Scale Simulation Random Variables (SSRVs), which can simplify the specification and inference of Bayesian PIMs. We show that SSRVs allow researchers to represent more realistic beliefs about scale than is possible with current methods.
2 Motivating Example
The analysis of gut microbiota data often centers around a $D\times N$ sequence count table $\mathbf{Y}$ with elements $\mathbf{Y}_{dn}$ denoting the number of sequenced (e.g., measured) DNA molecules that map to a particular microbial taxon $d$ in fecal sample $n$. The measurement process is complex and requires many experimental and computational steps to transform fecal samples into a count table $\mathbf{Y}$ (Caporaso et al.,, 2011). Researchers frequently analyze these data in the hopes of identifying which, if any, taxa change in abundance between cases and controls. This is referred to as differential abundance analysis. As a motivating example, we will discuss the complexity of this problem and later show that standard methods for performing differential abundance analysis can be very inaccurate.
One can consider the measurement process producing the sequence count table $\mathbf{Y}$ as a type of poll: there is a pool of microbes that could be measured, but only an arbitrary sized subset of those are actually collected (Fernandes et al.,, 2014; Lovell et al.,, 2011). Let $\mathbf{W}$ denote a $D\times N$ matrix with elements $\mathbf{W}_{dn}$ representing the true amount of taxon $d$ in the gut from which fecal sample $n$ was obtained. Note that this matrix can be uniquely described in terms of its composition $\mathbf{W}^{{\tiny\mbox{comp}}}$ and scale $W^{{\tiny\mbox{tot}}}$ via:
$\mathbf{W}_{dn}$ $=\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}\,\,W^{{\tiny\mbox{tot}}}_{n}$
$W^{{\tiny\mbox{tot}}}_{n}$ $=\sum_{d=1}^{D}\mathbf{W}_{dn}.$
This implies that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is a $D\times N$ matrix with columns that sum to $1$, and $W^{{\tiny\mbox{tot}}}$ is a positive-valued $N$-vector. When we say that the scale of the data is arbitrary, we mean that, even asymptotically, $\mathbf{Y}$ does not contain sufficient information to uniquely identify the system scale $W^{{\tiny\mbox{tot}}}$. This statement is formally defined in Section 4.
One of the most common analyses of microbiome data is called Differential Abundance Analysis (DA). In DA, there is a binary covariate ($x_{n}\in\{0,1\}$) associated with each measurement which denotes one of two mutually exclusive conditions (e.g., case versus control). The simplest estimand for DA is the log-fold-change (LFC). The LFC of taxon $d$, $\theta_{d}$, is often defined as the difference in means of log-transformed abundances (Love et al.,, 2014):
Estimating LFCs requires knowledge of the system scale $W^{{\tiny\mbox{tot}}}$. Based on the decomposition of $\mathbf{W}_{dn}$, we can write $\log\mathbf{W}_{dn}=\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}+\log W^{{\tiny\mbox{tot}}}_{n}$, thus
$=\theta^{\tiny\mbox{comp}}_{d}+\theta^{\tiny\mbox{tot}}.$ (2)
The main obstacle in obtaining accurate estimates of this estimand is that the scale of the data $\mathbf{Y}$ is typically arbitrary, and, as a result, we lack information about $W^{{\tiny\mbox{tot}}}$ and, therefore, $\theta^{\tiny\mbox{tot}}$.
To illustrate the problem, suppose that a researcher concludes that taxon $d$ is increased in cases as opposed to controls ($\theta_{d}>0$) based on the observation that the proportional abundance of taxon $d$ increases ($\theta^{\tiny\mbox{comp}}_{d}>0$). This conclusion is not supported by the observed data as knowledge of $\theta^{\tiny\mbox{comp}}_{d}$ is insufficient to determine $\theta_{d}$. Suppose the mean scale in the case condition is lower than the mean scale in the control condition ($\theta^{\tiny\mbox{tot}}<0$). Then, $\theta^{\tiny\mbox{comp}}_{d}$ will be spuriously large. In fact, if $|\theta^{\tiny\mbox{tot}}|>|\theta^{\tiny\mbox{comp}}_{d}|$, then $\theta^{\tiny\mbox{comp}}_{d}$ suggests that taxon $d$ is increasing when it is actually decreasing.
In Supplementary Section A, we provide a toy example with just $D=4$ taxa so that readers can walk through calculations themselves and see the spurious conclusions that can arise when analyzing data with arbitrary scale. Later, in Section 7, we provide a larger ($D=20$) simulation and an analysis of real data which demonstrates how two popular methods for DA analysis can demonstrate elevated Type I and/or Type-II error rates due to the problem of scale. For example, in Section 7.1, we demonstrate that two popular methods used for DA analysis have Type-I Error rates of over 70%. Contrary to standard statistical wisdom, we find these error rates actually increase with more data. Still, in those sections, we show that rigorous analyses are possible when statistical methods account for potential errors in modeling assumptions about scale.
3 Notation
@@
We use $P_{a,b}$ to denote a probability model which is a function mapping from a parameter space $\mathcal{A}\times\mathcal{B}$ to a set of distributions $\mathcal{P}$. For a random variable $x$ which can be described by this model, we use the notation $P_{a,b}(x)$ to denote specific distribution of $x$ within the set $\mathcal{P}$ (i.e., a point within the image of the model $P_{a,b}$). The identification region ${\mbox{Id}}_{a,b}(P)$ is the pre-image of the model at a specific distribution $P\in\mathcal{P}$:
@@
The marginal identification region of $b$ (${\mbox{Id}}_{b}(P)$) is defined analogously. Statements about the identifiability of a single parameter within a multi-parameter model refer to the marginal identification region: i.e., $a$ is identified at $P$ if $Id_{a}(P)$ is a singleton set. While identification can depend on the specific point $P\in\mathcal{P}$, for brevity, we will often omit this detail when the specific point can be inferred from context.
4 Formal Problem Statement
Here, we formally define what it means to ask a scale reliant question using data with arbitrary scale. As mentioned, we call this the problem of Scale Reliant Inference (SRI). First, we define the problem formally and then discuss how the problem statement relates to real-world problems such as DA analysis.
Consider the observed data $\mathbf{Y}$ as generated by a probability model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$. We say that $\mathbf{Y}$ has arbitrary scale if the parameter $W^{{\tiny\mbox{tot}}}$ is unidentified at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$. For simplicity, this article assumes that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is identified. We make this assumption to focus our attention on the impact of scale and not other sources of measurement limitations associated with $\mathbf{W}^{{\tiny\mbox{comp}}}$ (e.g., PCR bias (Silverman et al.,, 2021)). We will denote data generation models where $\mathbf{W}^{{\tiny\mbox{comp}}}$ can be identified but $W^{{\tiny\mbox{tot}}}$ cannot as data with arbitrary scale or compositional surveys. We choose this name to highlight the connection to compositional data. However, the setting we consider is more general than the classic CoDA setting. We do not impose the simplex constraint of compositional data, and, therefore, we can define and model compositional surveys where each measurement, $\mathbf{Y}_{\cdot n}$, could be integer-valued and need not sum to a constant value.
The research questions we consider can be framed as learning the map $\mathbf{Y}\mapsto\theta$ where $\theta:=\theta\left(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}\right)$. We call $\theta$ a target estimand because it is the quantity a researcher wants to estimate. We also introduce a nuisance function $\lambda=\lambda(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$ to ensure that the following concatenated function $f$ defined as
@@ 4 Formal Problem Statement
is invertible. With this condition, the model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$ can be reparameterized in terms of $(\theta,\lambda)$ so $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}=P_{f^{-1}(\theta,\lambda)}$. We define SRI in terms of the identifiability of $\theta$ within this reparameterized model.
Definition 1 (Scale Reliant and Scale Invariant Inference).
@@
To help motivate this definition, consider that, if $\mathbf{Y}$ is a compositional survey, the only way that $\theta$ can be unidentified is if it depends on $W^{{\tiny\mbox{tot}}}$ (since $\mathbf{Y}$ provides information about $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta$ is a known function). Therefore, our problem statement provides focus to our study: it restricts our attention to the setting where the lack of information about $W^{{\tiny\mbox{tot}}}$ contained in $\mathbf{Y}$ impedes our ability to uniquely identify $\theta$.
4.1 A Continuum between Scale Invariant and Scale Reliant Inference

### v5

@@ Abstract
Many scientific fields, including human gut microbiome science, collect multivariate count data where the sum of the counts is unrelated to the scale of the underlying system being measured (e.g., total microbial load in a subject’s colon). This disconnect complicates downstream analyses such as differential analysis in case-control studies. This article is motivated by a novel study of in vitro human gut microbiome models. Popular tools for analyzing these data led to dramatically elevated rates of both false positives and false negatives. To understand those failures, we provide a formal problem statement that frames these challenges of scale in terms of the classical theory of identifiability. We call this the problem of Scale Reliant Inference (SRI). We use this formulation to prove fundamental limits on SRI in terms of criteria such as consistency and type-I error control. We show that the failures of existing methods stem from a fundamental failure to properly quantify uncertainty in the system scale. We demonstrate that a particular type of Bayesian model called a Bayesian Partially Identified Model (PIMs) can correctly quantify uncertainty in SRI. We introduce Scale Simulation Random Variables (SSRVs) as a flexible and efficient approach to specifying and inferring Bayesian PIMs. In the context of both real and simulated data, we find SSRVs drastically decrease type-I and type-II error rates.
Multivariate Count Data ,
Partially Identified Models,
Compositional Data Analysis,
Microbiome Data,
keywords:
[…]
, , , , , , and
@@ 1 Introduction
This article is motivated by statistical challenges that arise in many areas of science and have become a central focus of modern microbiome research. Microbiome sequence count data are typically represented as an $D\times N$ table of counts $\mathbf{Y}$ with elements $\mathbf{Y}_{dn}$ representing the number of DNA sequences observed that map to the $d$-th microbial taxon in the $n$-th biological sample (e.g., a fecal sample from the $n$-th study participant). The measurement process is non-trivial yet, for the purposes of the present article, can be thought of as a multivariate sample: there is a large pool of microbes in the human gut that could be measured, and our measurements capture a small, arbitrary-sized sample of that pool (Silverman et al., 2018; Roche and Mukherjee, 2022). By arbitrary sized, we allude to the fact that the size of the sample (the sequencing depth $\sum_{d}\mathbf{Y}_{dn}$) is typically independent of the scale of the biological system from which that sample was obtained (e.g., the total number of microbes in the subject’s colon) (Lovell et al., 2011; Props et al., 2017; Vandeputte et al., 2017). As a result, many authors call these data compositional, alluding to the idea that the data only provides information about the proportional abundance of each microbe and not their absolute abundance (Gloor et al., 2016a; Lovell et al., 2011; Gloor et al., 2017). The problem is that many of the scientific questions asked of these data require information about absolute abundances rather than just proportional abundances.
For concreteness, this article introduces a novel microbiome study that aimed to quantify the limits of in vitro artificial gut vessels as models of human gut microbiota. Eight artificial gut systems were inoculated with a human fecal microbial community sample. The inoculated vessels incubated those communities for two weeks. Our goal was to identify which of the $D$ microbial taxa in the community can grow within the vessels, quantified as an increased abundance on day 14 compared to day 1. This is an example of a differential abundance analysis which aims to identify taxa that differ in abundance between two experimental conditions. For clarity, we are not interested in changes in proportional abundances as, for example, a taxon may decrease in relative abundance between days 1 and 14 simply because its growth was outpaced by some other microbe that grew more rapidly. While differential abundance analyses are performed in many, if not most, microbiome studies (Lin and Peddada, 2020), we find that popular tools for the task lead to exceedingly high rates of false positives and false negatives in our study. In this article, we argue that this is not simply an issue of data size, measurement error, or choice of parametric model but a fundamental failure of these methods to properly quantify uncertainty originating from the lack of scale information in these data.
The main contributions of this article are the following. First, we provide a formal problem statement that defines what it means to ask a question that depends on scale using data that lacks information about scale. We call this the problem of Scale Reliant Inference (SRI) and frame it in terms of the classical theory of identifiability. We discuss the advantages of our formulation over the related problem of Compositional Data Analysis (CoDA). Second, we use this formulation to prove fundamental limits on SRI. Contrary to numerous published claims, we prove that classical inferential criteria such as consistency and Type-I error control cannot be attained in many SRI problems. In contrast, we show that Bayesian inferential criteria can be achieved by a specialized class of models called Bayesian Partially Identified Models (PIMs). We introduce a simulation framework called Scale Simulation Random Variables (SSRVs), which can simplify the specification and inference of Bayesian PIMs for SRI. Finally, we demonstrate SSRVs through both simulated and real data analysis; in particular, we use SSRVs to tackle the differential abundance analysis at the core of the artificial gut vessel study introduced above. In the context of real data, we find SSRVs control false discovery rates while popular alternatives display rates as high as 30 percent. Moreover, SSRVs drastically reduce false rates, identifying over three times more true positives than alternatives. Our methods have already been implemented into popular software tools (Nixon, Gloor and Silverman, 2024). Moreover, Our findings, namely the dramatic decreases in Type-I and/or Type-II error rates compared to existing tools, has been validated in the context of other microbiome datasets (Nixon, Gloor and Silverman, 2024) and meta-transcriptomics (Dos Santos et al., 2024).
While this article focuses on applications to microbiome analysis, the problems discussed and the theory and methods developed are more general. Beyond microbiome studies, scale reliant inference poses problems in other areas of biomedicine (e.g., gene expression studies) (Fernandes et al., 2014; Lovell et al., 2011), macro-ecology (Billheimer, Guttorp and Fagan, 2001; Jackson, 1997; Murphy et al., 2016), political poling (Topîrceanu and Precup, 2020), demography (Rossi et al., 1987), and epidemiology (Silverman, Hupert and Washburne, 2020). In all these fields, multivariate count data is collected where the total counts are arbitrary and unrelated to the measured systems’ underlying scale (i.e., size).
2 A Theory for Scale Reliant Inference
In this section we introduce a theory for scale reliant inference from data that lacks information about scale. We also discuss implications of this theory on inference procedures. First we set notation that will be used throughout the paper.
2.1 Notation
@@
We use $P_{a,b}$ to denote a probability model which is a function mapping from a parameter space $\mathcal{A}\times\mathcal{B}$ to a set of distributions $\mathcal{P}$. For a random variable $x$ the notation $P_{a,b}(x)$ denotes a specific distribution of $x$ within the set $\mathcal{P}$, a point within the image of the model $P_{a,b}$. The identification region ${\mbox{Id}}_{a,b}(P)$ is the pre-image of the model at a specific distribution $P\in\mathcal{P}$:
@@
The marginal identification region of $b$ (${\mbox{Id}}_{b}(P)$) is defined analogously. Statements about the identifiability of a single parameter within a multi-parameter model refer to the marginal identification region: i.e., $a$ is identified at $P$ if ${\mbox{Id}}_{a}(P)$ is a singleton set. While identification can depend on the specific point $P\in\mathcal{P}$, for brevity, we will often omit this detail when the specific point can be inferred from context.
We consider a $D\times N$ matrix of observed data $\mathbf{Y}$ which is a measurement of a $D\times N$ dimensional underlying system $\mathbf{W}$. We place no mathematical restrictions on $\mathbf{Y}$ and only require that $\mathbf{W}$ is positive-valued (integer or continuous) such that the elements $\mathbf{W}_{dn}$ represent the amount of class $d$ in source $n$ (e.g., the abundance of taxon $d$ in subject $n$’s gut). The system $\mathbf{W}$ can be uniquely decomposed into its composition ($\mathbf{W}^{{\tiny\mbox{comp}}}$, proportions) and scale ($W^{{\tiny\mbox{tot}}}$, totals) via:
$\mathbf{W}_{dn}$ $=\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}\,\,W^{{\tiny\mbox{tot}}}_{n}$ (1)
$W^{{\tiny\mbox{tot}}}_{n}$ $=\sum_{d=1}^{D}\mathbf{W}_{dn}.$
This implies that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is a $D\times N$ matrix with columns that sum to $1$, and $W^{{\tiny\mbox{tot}}}$ is a positive-valued $N$-vector. We denote rows of a matrix with a placeholder dot, e.g., $\mathbf{W}_{\cdot n}$ is the $n$th row of $\mathbf{W}$.
2.2 Formal Problem Statement
This subsection presents a formal definition of Scale Reliant Inference (SRI) followed by an example of SRI in microbiome data analyses.
Assume the observed data $\mathbf{Y}$ are generated following a probability model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$. We say that $\mathbf{Y}$ lacks information about scale if the parameter $W^{{\tiny\mbox{tot}}}$ is unidentified at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$. For simplicity, this article assumes that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is identified. We make this assumption to focus our attention on the impact of scale and not other sources of measurement limitations associated with $\mathbf{W}^{{\tiny\mbox{comp}}}$ (such as PCR bias in microbiome studies Silverman et al. (2021)). We denote data generation models where $\mathbf{W}^{{\tiny\mbox{comp}}}$ can be identified but $W^{{\tiny\mbox{tot}}}$ cannot as compositional surveys. We choose this name to highlight the connection to compositional data. That connection, along with important distinctions between SRI and compositional data analysis are reviewed in Section 2.3.
SRI can be framed as learning the map $\mathbf{Y}\mapsto\theta$ where $\theta:=\theta\left(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}\right)\in\Theta$. We call $\theta$ a target estimand because it is the quantity one wants to estimate. We also introduce a nuisance function $\lambda=\lambda(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$ to ensure that the following concatenated function $f$ defined as
@@ 2.2 Formal Problem Statement
is invertible. With this condition, the model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$ can be reparameterized in terms of $(\theta,\lambda)$ such that $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}=P_{f^{-1}(\theta,\lambda)}$. We define SRI in terms of the identifiability of $\theta$ within this reparameterized model.
Definition 2.1 (Scale Reliant and Scale Invariant Inference).
@@
To help motivate this definition, consider that, if $\mathbf{Y}$ is a compositional survey, the only way that $\theta$ can be unidentified is if it depends on $W^{{\tiny\mbox{tot}}}$ (since $\mathbf{Y}$ provides information about $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta$ is a known function). Therefore, our problem statement provides focus to our study: it restricts attention to the setting where the lack of information about $W^{{\tiny\mbox{tot}}}$ contained in $\mathbf{Y}$ impedes our ability to uniquely identify $\theta$.
Example (Differential Abundance Analysis from Microbiome Data).
The analysis of human gut microbiome data often centers around a $D\times N$ sequence count table $\mathbf{Y}$ with elements $\mathbf{Y}_{dn}$ denoting the number of sequenced (measured) DNA molecules that map to a particular microbial taxon $d$ in fecal sample $n$. For concreteness, we provide a toy dataset ($D=4$) in Supplementary Section A. The measurement process is complex and requires many experimental and computational steps to transform fecal samples into a count table $\mathbf{Y}$ (Caporaso et al., 2011). We model the measurement process as random, there exist some true system with composition $\mathbf{W}^{{\tiny\mbox{comp}}}$ and scale $W^{{\tiny\mbox{tot}}}$ but it is only observed through the noisy measurements $\mathbf{Y}$. While the true number of microbes in a system is integer valued; for simplicity, we assume that those counts are large enough that we can approximate $W_{dn}\in\mathbb{R^{+}}$.
Many authors call these data compositional owing to the measurement process. The measurement process can be thought of as a type of poll: there is a pool of microbes that could be measured, but only an arbitrarily sized subset of those are (Fernandes et al., 2014; Lovell et al., 2011). Moreover, due to complexities such as DNA amplification, it is not even possible to assume that the system has more sequences than the sample ($\mathbf{W}_{dn}>\mathbf{Y}_{dn}$). Overall, the scale of the underlying system is lost through this process leading to data that provides information about the system composition but not the system scale (Vandeputte et al., 2017; Props et al., 2017). We model sequence count data as a compositional survey to reflect this feature of the data.
The goal of many microbiome studies is to determine which, if any, taxa differ in abundance between two mutually exclusive conditions (e.g., disease versus health). Letting $x_{n}\in\{0,1\}$ denote the condition of sample $n$, these differential abundance analyses, often center on estimation of Log-Fold-Changes (LFCs):
LFCs are therefore the target estimand in differential abundance analyses. Combining Equations (1) and (2) leads to the following relationship
$=\theta^{\tiny\mbox{comp}}_{d}+\theta^{\tiny\mbox{tot}}.$ (3)
In words, the LFC estimand requires knowledge of how the proportional abundance of the $d$-th taxon changes between conditions ($\theta^{\tiny\mbox{comp}}_{d}$) as well as how total microbial load changes between conditions ($\theta^{\tiny\mbox{tot}}$). Because microbiome data ($\mathbf{Y}$) is a compositional survey, it follows that $\theta^{\tiny\mbox{tot}}$, and therefore $\theta_{d}$, are not identifiable. By definition, differential abundance analysis from microbiome data is a problem of Scale Reliant Inference.
Before progressing, we discuss a subtlety of our model that is a source of frequent confusion. There are often two sources of uncertainty in SRI. Using LFC estimation from microbiome data as an example, there will will be uncertainty in $W^{{\tiny\mbox{tot}}}$, and therefore $\theta^{\tiny\mbox{tot}}$, since both are unidentified. Yet we will also have uncertainty in $\mathbf{W}^{{\tiny\mbox{comp}}}$, and therefore $\theta^{\tiny\mbox{comp}}$, since we only observe a noisy, finite sample $\mathbf{Y}$ and not $\mathbf{W}^{{\tiny\mbox{comp}}}$ directly. To clarify, our assumption that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is identifiable does not imply that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is known. Identifiability is a feature of the distribution of observables $P_{W^{{\tiny\mbox{tot}}},\mathbf{W}^{{\tiny\mbox{comp}}}}(\mathbf{Y})$; yet, we cannot know that distribution without potentially infinite number of replicate studies. In microbiome data analyses, we can have substantial uncertainty in $\mathbf{W}^{{\tiny\mbox{comp}}}$ due to the variability in the count data $\mathbf{Y}$. Still, this article places comparatively less emphasis on this later source of compositional uncertainty as there is already substantial literature on that subject which we will use in later sections (e.g., Gloor et al. (2016b); Silverman et al. (2018)).
2.3 Comparing Compositional Data Analysis and SRI
This article does not require a detailed understanding of Compositional Data Analysis (CoDA). Still, CoDA has become increasingly popular for analyzing human gut microbiome data, and as a result, we provide a brief comparison for readers who are already familiar with that literature. A broader review of CoDA and a more detailed introduction to CoDA in the context of microbiome data can be found in Greenacre et al. (2023) and Gloor et al. (2016a), respectively.
In brief, CoDA studies simplex-valued data: multivariate, continuous, positive-valued vectors that sum to a constant value. In the context of microbiome data, researchers often identify each datum $\mathbf{Y}_{\cdot n}$ with a single vector on simplex via normalization (e.g., $\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}=\mathbf{Y}_{\cdot n}/\sum_{d}Y_{dn}$(Srinivasan, Xue and Zhan, 2020). While we have used our notation $\mathbf{W}^{{\tiny\mbox{comp}}}$, it is important to note that CoDA is not formulated with any explicit notion of the system being measured. Said differently, CoDA does not distinguish between the observed data and the underlying system being studied. This is a critical difference between SRI and CoDA. We use the system $\mathbf{W}$ as a conceptual tool to define the true value of our estimand $\theta(\mathbf{W})$. This, combined with our probabilistic definition of compositional surveys, will allow us to investigate SRI and motivate our methods using classical inferential criteria (e.g., consistency and type-I error control). Lacking these structures, CoDA relies on three non-standard axioms to motivate its approach (Pawlowsky-Glahn, Egozcue and Tolosana-Delgado, 2015, Section 2.2). Chief among these is the principle of scale invariance, which states that any analysis of compositional data should be invariant to the arbitrary scale of the data. As a result of this axiom, CoDA restricts its attention to log-ratio estimands which are scale invariant: e.g., estimands of the form $\gamma=f(\eta_{1},\dots,\eta_{N})$ where $\eta_{n}=\log(\mathbf{W}_{in}/\mathbf{W}_{jn})$ (Lovell, Chua and McGrath, 2020). In contrast, SRI studies the estimation of scale reliant estimands.
Despite these critical differences, we choose terminology that is consistent with CoDA. For example, we use the term scale to refer to the sums $W^{{\tiny\mbox{tot}}}_{n}=\sum_{d}\mathbf{W}_{dn}$ of the system in the same way that CoDA uses the term to refer to the sums of the data $\sum_{d}\mathbf{Y}_{dn}$. We choose the term compositional survey to distinguish our probability model from the mathematical definition of compositional data while highlighting the intuitive relationship between them. Note our definition of compositional surveys does not require that the measurements $\mathbf{Y}_{\cdot n}$ are continuous or sum to a constant value. This is appealing as microbiome data lack information about the system scale yet are count vectors that do not sum to a constant value.
The implications of these differences between CoDA and SRI are discussed in later sections and summarized in Section 5.
2.4 The Continuum between SII and SRI